OpenStudy (anonymous):

3* Cube root of (x^2)- 10* Cube root of Radical X-9=o

OpenStudy (istim):

And you're solving for x?

OpenStudy (anonymous):



OpenStudy (anonymous):

and i didnt mean radical x, haha, its already under a cube root

OpenStudy (istim):

I'm afk for a while. Sorry. Someone will more likely be here to help soon.

OpenStudy (anonymous):

let u=cubert(x) 3u^2-10u-8=0 (3u+2)(u-4)=0 so 3u+2=0 3*cubert(x)=-2 x=(-2/3)^3 x=-8/27 cubert(x)=4 x=4^3 x=64